OpenStudy (anonymous):

Dayton is due north of Richmond, and Easton is 16 mi due east of Richmond. Dayton is 32 mi from Easton. How far apart are Dayton and Richmond? Round to the nearest tenth of a mile.

OpenStudy (godlovesme):

\(\huge\color{purple}{WELCOME~TO~OS~!!!~:)}\)

OpenStudy (alexandervonhumboldt2):



OpenStudy (alexandervonhumboldt2):

by pythogorian theorem 16^2 + x^2=32^2

OpenStudy (alexandervonhumboldt2):

x^2=32^2 -16^2 just do some simple math

OpenStudy (anonymous):

would it be 35.77

OpenStudy (alexandervonhumboldt2):

nope

OpenStudy (anonymous):

or 27.72

OpenStudy (alexandervonhumboldt2):

this is correct 27.72 rounded

OpenStudy (anonymous):

yay thx

OpenStudy (anonymous):

what was it